The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 44 May 15, 2000

Contents

Abstracts of 13 refereed papers  1
Abstracts of 3 conference contributions 10

Abstracts of Refereed Papers

Formation of Cavities, Filaments, and Clumps by the Non-linear Development of Thermal and Gravitational Instabilities

in the Interstellar Medium under Stellar Feedback

K. Wada^{1}, M. Spaans^{2}, and S. Kim^{2}

1 National Astronomical Observatory, Mitaka,181-8588, Japan 2 Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA 3 Astronomy Department, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA

Based on our high resolution, two-dimensional hydrodynamical simulations, we propose that large cavities may be formed by the nonlinear development of the combined thermal and gravitational instabilities, without need for stellar energy injection in a galaxy modeling the Large Magellanic Clouds (LMC). Our numerical model of the star formation allows us to follow the evolution of the blastwaves due to supernovae in the inhomogeneous, multi-phase, and turbulent-like media self-consistently. Formation of kpc-scale inhomogeneity, such as cavities, observed H i map of the LMC, is suppressed by frequent supernovae (average supernova rate for the whole disk is ∼ 0.001 yr^{−1}). However the supernova explosions are necessary for the hot component (T_{g} > 10^{6−7} K). Position-velocity maps show that kpc-scale shells/arcs formed through the nonlinear evolution in a model without stellar energy feedback has similar kinematics to explosional phenomena, such as supernovae. We also find that dense clumps and filamentary structure are formed due to a natural consequence of the non-linear evolution of the multi-phase ISM. Although the ISM in a small scale looks turbulent-like and transient, the global structure of the ISM is quasi-stable. In the quasi-stable phase, the volume filling factor of the hot, warm, cold components are ∼ 0.2, ∼ 0.6, and ∼ 0.2, respectively. We compare the observations of H i and molecular gas of the LMC with the numerically obtained H i and CO brightness temperature distribution. The morphology and statistical properties of the numerical H i and CO maps are discussed. We find that the cloud mass spectrum of our models represent a power-law shape, but their slopes change between models with and without the stellar energy injection, and also the slope depends on the threshold brightness temperature of CO.

Accepted by: The Astrophysical Journal (to appear in Vol. 540)

For preprints, contact wada.keiichi@nao.ac.jp
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0005330

Interpretation of the Spatial Power Spectra of Neutral Hydrogen in the Galaxy and in the Small Magellanic Cloud

Itzhak Goldman

School of Physics and Astronomy, Sackler Faculty of Exact Sciences, Tel Aviv University, Tel Aviv 69978, Israel

Recent 21 cm radio observations of H i regions in the Small Magellanic Cloud have revealed spatial power spectra of the intensity, which are quite similar in shape to those previously deduced for the Galaxy. The similarity, in spite the differences in the physical parameters between the Galaxy and the SMC, suggests that the shape of the power spectra reflects some underlying mechanism which is not too sensitive to the environmental specifics. In this paper we present an interpretation for the observational power spectra in terms of a large scale turbulence in the interstellar medium, in which the emitting H i regions are embedded. The turbulence gives rise to density fluctuations which lead to the observed intensity fluctuations, in the H i regions. The observational power spectra are used to deduce the turbulence spectral function.

In the SMC, the turbulence largest eddies are comparable in scale to the SMC itself. This implies that turbulent mixing should have smoothed out any large scale abundance gradients. Indeed, this seems to be the case, observationally.

The turbulence is also expected to amplify and shape up the large scale magnetic field. Indeed, the observational data indicate the existence of a large scale disordered field of the strength expected from energy equilibrium with the turbulent velocity field. The large scale turbulence is most probably generated by instabilities in the large scale flows induced by the tidal close encounter with the LMC ∼ 2 × 10^{8}yr ago. The life-time of the largest eddies is ∼ 4 × 10^{8}yr so the turbulence had not yet enough time to decay and persists even though the energy source is no longer there.

Accepted by: The Astrophysical Journal

For preprints, contact goldman@wise1.tau.ac.il
Also available from the URL http://arXiv.org/abs/astro-ph/0005023

H_{2} measurements in the Magellanic Clouds

Philipp Richter^{1,2}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Department of Astronomy, University of Wisconsin-Madison, Madison, WI 53706, USA

More than 20 years after the Copernicus satellite, ORFEUS allows the investigation of molecular hydrogen (H_{2}) in the diffuse interstellar medium by way of FUV absorption spectroscopy once again. This time, targets in the Magellanic Clouds were also observed, allowing the first investigation of cold H_{2} in a metal-poor environment outside the Milky Way disk. This paper presents new H_{2} measurements in LMC gas along the lines of sight toward HD 269698 (Sk −67 166), HD 269546 (Sk −68 82) and HD 36402 (Sk −67 104) and summarizes the ORFEUS H_{2} measurements in the the Magellanic Clouds. For seven lines of sight we investigate correlations between N(H i), N(H_{2}), E(B − V) and compare the results with H_{2} observations in the Milky Way disk. We find that the abundance of H_{2} is low in comparison to the total gas quantity and speculate that the fraction of hydrogen in molecular form is limited by the lower dust content of the Magellanic Clouds.

Accepted by: Astronomy & Astrophysics, Main Journal

For preprints, contact richter@astro.wisc.edu

Multiwavelength observations of N 66 in the SMC: unveiling photodissociation interfaces and star formation

M. Rubio^{1} A. Contursi ^{2,3,4} J. Lequeux ^{4} R. Probst ^{5} R. Barbá ^{6} F. Boulanger^{6} D.

Cesarsky ^{7,9} and R. Maoli^{8}

1 Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile 2 IPAC, Caltech, MS 100–22, Pasadena, CA 91125, USA 3 SAp/DAPNIA/DSM, CEA-Saclay, F-91191 Gif sur Yvette CEDEX, France 4 DEMIRM, Observatoire de Paris, 61 Avenue de l’Observatoire, F-75014 Paris, France 5 Cerro Tololo Inter-American Observatory, NOAO, Casilla 603, La Serena, Chile 6 Facultad de Ciencias Astronómicas y Geofisícas, Universidad Nacional de La Plata, Paseo del Bosque s/n, 1900 La Plata, Argentina 7 Institut d’Astrophysique Spatiale, Bat. 121, Université Paris XI, 91450 Orsay CEDEX, France 8 Institut d’Astrophysique, 98 bis Boulevard Arago, 75014 Paris, France 9 Max-Planck-Institut für Extraterrestrische Physik, P.O. Box 1603, D-85740 Garching, Germany

We present new observations of the major star–forming region N 66 in the Small Magellanic Cloud and of its surroundings, which add to those presented in Contursi et al. 2000 (Contursi A., Lequeux J., Cesarsky D. et al. 2000, A&A, submitted, Paper I). High–sensitivity CO observations allowed the detection of molecular gas associated with the H ii region, for which a high–resolution image in [O iii] λ 5007 is presented. We also present images in the v(1-0) S(1) line of H_{2} at 2.12 μm and in the adjacent continuum. This material reveals an interesting photodissociation region. We show that the molecular gas that has not yet been photodissociated by the UV radiation of the stars is in hot, dense clumps with a very small surface filling factor. We discovered several embedded stars or stellar associations, and suggest that three successive stellar generations have already taken place in less than 3 million years.

Accepted by: Astronomy & Astrophysics Main Journal

For preprints, contact mrubio@das.uchile.cl

On the Emergence and Discovery of Hot Spots in Supernova Remnant 1987A

Stephen S. Lawrence^{1}, Ben E. Sugerman^{1}, Patrice Bouchet^{2}, Arlin P. S. Crotts^{1}, Robert Uglesich^{1}, and Steve Heathcote^{2}

1 Dept. of Astronomy, Columbia University, New York, NY 1002 7, USA 2 Cerro Tololo Inter-American Observatory, La Serena, Chile

We present the discovery of several new regions of interaction between the ejecta and equatorial circumstellar ring of SN1987A, an interaction leading to a much expanded development of the supernova remnant. We also trace the development of the first such “hot spot,” discovered in 1997, back to 1995. Later hot spots seem to have emerged by early 1999. We discuss mechanisms for the long delay between the first and later spots.

Submitted to: Astrophysical Journal Letters

For preprints, contact lawrence@astro.columbia.edu
Also available from the URL http://arXiv.org/abs/astro-ph/0004191

Far Ultraviolet Spectroscopic Explorer Observations

of the Supernova Remnant N49 in the Large Magellanic Cloud

William P. Blair^{1}, Ravi Sankrit^{1}, R.L. Shelton^{1}, Kenneth R. Sembach^{1}, H. Warren Moos^{1}, John C. Raymond^{2}, Donald G. York^{3}, Paul D. Feldman^{1}, Pierre Chayer^{1}, Edward M. Murphy^{1}, David J. Sahnow^{1}, and Erik Wilkinson^{4}

1 Department of Physics & Astronomy, The Johns Hopkins University, 3400 N. Charles St., Baltimore, MD 21218, USA 2 Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA 3 University of Chicago, 5640 S. Ellis Ave., Chicago, IL 60637, USA 4 University of Colorado, Campus Box 593, 1255 38th St., Boulder, CO 80302, USA

We report a Far Ultraviolet Spectroscopic Explorer satellite observation of the supernova remnant N49 in the Large Magellanic Cloud, covering the 905 – 1187 Å spectral region. A 30′′ square aperture was used, resulting in a velocity resolution of ∼100 km s^{−1}. The purpose of the observation was to examine several bright emission lines expected from earlier work and to demonstrate diffuse source sensitivity by searching for faint lines never seen previously in extragalactic supernova remnant UV spectra. Both goals were accomplished. Strong emission lines of O vi λλ1031.9, 1037.6 and C iii λ977.0 were seen, Doppler broadened to ±225 km s^{−1} and with centroids red-shifted to 350 km s^{−1}, consistent with the LMC. Superimposed on the emission lines are absorptions by C iii and O vi λ1031.9 at +260 km s^{−1}, which are attributed to warm and hot gas (respectively) in the LMC. The O vi λ1037.6 line is more severely affected by overlying interstellar and H_{2} absorption from both the LMC and our galaxy. N iii λ989.8 is not seen, but models indicate overlying absorption severely attenuates this line. A number of faint lines from hot gas have also been detected, many of which have never been seen in an extragalactic supernova remnant spectrum.

Accepted by: The Astrophysical Journal (Letters); will appear in a special issue devoted to the first scientific results from the FUSE mission For preprints, contact wpb@pha.jhu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0005092

Measuring the ionization of O star winds

Luciana Bianchi^{1,2,3}, J. B. Hutchings^{1,4}, Alex W. Fullerton^{5,2}, Donald C. Morton^{1,4}, Raymond G. Ohl^{2}, George Sonneborn^{6}, Derck Massa^{7}, K.R. Brownsberger^{8}, Allan J.

Willis^{9}

1 Observer with the NASA-CNES-CSA Far Ultraviolet Spectroscopic Explorer (FUSE) which is operated by Johns Hopkins University, under NASA contract NAS5-32985 2 Center for Astrophysical Sciences, The Johns Hopkins University, Dept. of Physics and Astronomy, 239 Bloomberg Center, 3400 N. Charles St., Baltimore, MD 21218, USA 3 Osservatorio Astronomico di Torino, I-10025 Pino Torinese (TO), Italy 4 Herzberg Inst. of Astrophysics, National Research Council of Canada, 5071 W. Saanich Rd., Victoria, B.C., Canada, V8X 4M6 5 Dept. of Physics & Astronomy, University of Victoria, P.O. Box 3055, Victoria, BC, V8W 3P6, Canada 6 Lab. for Astro. and Solar Physics, Code 681.0, NASA GSFC, Greenbelt, MD 20771, USA 7 Raytheon ITSS, Code 681.0, NASA GSFC, Greenbelt, MD 20771, USA 8 Center for Astrophysics and Space Astronomy, University of Colorado, Boulder, CO 80309, USA 9 Dept. of Physics and Astronomy, University College London, Gower St. London, UK

We present an analysis of wind line profiles from FUSE spectra of two O7 supergiants in the Large and Small Magellanic Clouds (Sk −67^{∘}111 and AV 232 respectively). Model fits yield the column densities of S iv, S vi, P iv, P v, N iii, and N iv, providing the first direct measurement of the ionization balance in stellar winds. The ratios of S iv/S vi and P iv/P v are consistently lower in the LMC star. IUE and HST archival spectra are also used to measure N iv and N v, but the much higher optical depth makes the N v measurements inconclusive. The velocity and optical depth distributions in the wind are qualitatively similar between the two stars, when scaled to their terminal velocities. The terminal velocities are different, with AV 232 being lower (as found previously in SMC stars and linked to lower metallicity). These are the first results from a program to investigate wind ionisation and velocity structure among hot stars in local galaxies, and demonstrate the higher accuracy in measuring column densities of less abundant ions such as phosphorus and sulfur, observable in the FUSE range.

Accepted by: The Astrophysical Journal Letters; will appear in a special issue devoted to the first scientific results from the FUSE mission For preprints, contact bianchi@stsci.edu

FUSE Observations of the Stellar Winds of Two O7 Supergiants in the Magellanic Clouds

A.W. Fullerton^{1,2}, P.A. Crowther^{3}, O. De Marco^{3}, J.B. Hutchings^{4}, L. Bianchi^{2,5}, K.R.

Brownsberger^{6}, D.L. Massa^{7}, D.C. Morton^{4}, B.L. Rachford^{6}, T.P. Snow^{6}, G. Sonneborn^{8}, J. Tumlinson^{6}, and A.J. Willis^{3}

1 Dept. of Physics & Astronomy, University of Victoria, P.O. Box 3055, Victoria, BC, V8W 3P6, Canada 2 Dept. of Physics & Astronomy, The Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21286, USA 3 Dept. of Physics & Astronomy, University College London, Gower Street, London WC1E 6BT, England 4 Herzberg Institute of Astrophysics, NRC of Canada, 5071 West Saanich Road, Victoria, BC V8X 4M6, Canada 5 Osservatorio Astronomico di Torino, I-10025 Pino Torinese (TO), Italy 6 CASA, University of Colorado at Boulder, Campus Box 389, Boulder, CO 80309, USA 7 Raytheon ITSS, NASA’s Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA 8 LASP, NASA’s Goddard Space Flight Center, Code 681, Greenbelt, MD 20771, USA

We compare the stellar wind features in far-UV spectra of Sk −67^{∘}111, an O7 Ib(f) star in the LMC, with Sk 80, an O7 Iaf+ star in the SMC. The most striking differences are that Sk 80 has a substantially lower terminal velocity, much weaker O vi absorption, and stronger S iv emission. We have used line-blanketed, hydrodynamic, non-LTE atmospheric models to explore the origin of these differences. The far-UV spectra require systematically lower stellar temperatures than previous determinations for O7 supergiants derived from plane-parallel, hydrostatic models of photospheric line profiles. At these temperatures, the O vi in Sk −67^{∘}111 must be due primarily to shocks in the wind.

Accepted by: Astrophysical Journal Letters; will appear in a special issue devoted to the first scientific results from the FUSE mission. For preprints, contact awf@pha.jhu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0005039

Far-UV FUSE spectroscopy of the O vi resonance doublet in Sand 2 (WO)

Paul A. Crowther^{1}, A.W. Fullerton^{2,3}, D.J. Hillier^{4}, K. Brownsberger^{5}, L. Dessart^{1,6},

A.J. Willis^{1}, O. De Marco^{1}, M.J. Barlow^{1}, J.B. Hutchings^{7}, D.L. Massa^{8},

D.C. Morton^{7} & G. Sonneborn^{9}

(1) Dept. of Physics & Astronomy, UCL, Gower St, London WC1E 6BT, U.K.

(2) Dept. of Physics & Astronomy, University of Victoria, P.O. Box 3055, Victoria, BC, V8W 3P6, Canada.

(3) Dept. of Physics & Astronomy, Johns Hopkins University, 3400 North Charles St., Baltimore, MD 21218, USA

(4) Dept. of Physics & Astronomy, University of Pittsburgh, 3941 O’Hara Street, PA 15260, USA

(5) Center for Astrophysics & Space Astronomy, Campus Box 389, Boulder, CO 80309, USA

(6) Present Address: Département de Physique, Université Laval, Québec, QC, Canada, G1K 7P4

(7) D.A.O., Herzberg Institute of Astrophysics, 5071 W. Saanich Road, Victoria BC, Canada, V8X 4M6

(8) Raytheon ITSS, NASA’s Goddard Space Flight Center, Code 631, Greenbelt, MD 20771, USA

(9) Laboratory for Astronomy and Solar Physics, NASA/GSFC, Code 681, Greenbelt, MD 20771, USA

We present Far-Ultraviolet Spectroscopic Explorer (FUSE) spectroscopy of Sand 2, a LMC WOtype Wolf-Rayet star, revealing the O vi resonance P Cygni doublet at 1032–38Å. These data are combined with HST/FOS ultraviolet and Mt Stromlo 2.3m optical spectroscopy, and analysed using a spherical, non-LTE, line-blanketed code. Our study reveals exceptional stellar parameters: T_{∗} ∼150,000K, v_{∞}=4100 km s^{−1}, log (L/L_{⊙})=5.3, and Ṁ = 1 × 10^{−5}M_{⊙}yr^{−1}, if we adopt a volume filling factor of 10%. Elemental abundances of C/He∼ 0.7 ± 0.2 and O/He∼ 0.15_{−0.05}^{+0.10} by number qualitatively support previous recombination line studies. We confirm that Sand 2 is more chemically enriched in carbon than LMC WC stars, and is expected to undergo a supernova explosion within the next 5×10^{4}yr.

To appear in: Astrophysical Journal Letters; will appear in a special issue devoted to the first scientific results from the FUSE mission. For preprints, contact pac@star.ucl.ac.uk

Also available from the URL http://www.star.ucl.ac.uk/∼pac/publications.html

The He II Emitting Nebula N44C in the LMC: Optical/UV Spectroscopy of the Nebula and its Ionizing Star

D.R. Garnett^{1}, V.C. Galarza,^{2}, and Y.-H. Chu^{3}

1 Steward Observatory, University of Arizona, 933 N. Cherry Ave., Tucson, AZ 85721, USA 2 Astronomy Department, New Mexico State University, Las Cruces, NM 88003-8001, USA 3 Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA

We present HST spectroscopy and imaging, along with new ground-based spectroscopy and ROSAT HRI imaging, of the He ii emitting nebula N44C and its ionizing star. A GHRS spectrum of the ionizing star yields a spectral type of about O7 for the star. The lack of P Cygni profiles for Si iv and C iv indicates that the star is not a supergiant. The nebular abundances in the ionized gas are consistent with average abundances for LMC H ii regions, with the possible exception that nitrogen may be enhanced. Enrichment by a former evolved companion star is not evident. A long-slit echelle spectrogram in Hα + [N ii] shows no evidence for high-velocity gas in N44C. This rules out high-velocity shocks as the source of the nebular He ii emission. A 108 ks ROSAT HRI image of N44C shows no X-ray point source to a 3σ upper limit L_{X} < 10^{34} erg s^{−1} in the 0.1-2.0 keV band. Based on new measurements of the electron density in the He ii emitting region, we derive recombination timescales of ≈ 20 yrs for He^{+2} and ≈ 4 yrs for Ne^{+4}. If N44C is a fossil X-ray ionized nebula, this places severe constraints on when the putative X-ray source could have turned off. The presence of strong [Ne iv] emission in the nebula is puzzling if the ionizing source has turned off. It is possible the system is related to the Be X-ray binaries, although the O star in N44C does not show Be characteristics at the present time. Monitoring of X-rays and He ii emission from the nebula, as well as a radial velocity study of the ionizing star, are needed to fully understand the emission line spectrum of N44C.

Submitted to: The Astrophysical Journal

For preprints, contact dgarnett@as.arizona.edu
Also available from the URL http://ocotillo.as.arizona.edu/~dgarnett/Papers/N44pp.ps

Doubling the number of Be/X-ray binaries in the SMC

F. Haberl and M. Sasaki

Max-Planck-Institut für extraterrestrische Physik, Garching, Germany

A correlation of X-ray source and Halpha emission-line object catalogues in the Small Magellanic Cloud (SMC) shows that more than two thirds of the optically identified Be stars in Be/X-ray binaries are found as emission-line objects in the catalogues. On the basis of this result we propose up to 25 X-ray sources mainly from recent ROSAT catalogues as new Be/X-ray binaries and give their likely optical counterparts. Also for the five yet unidentified X-ray pulsars in the SMC we propose emissionline stars as counterparts. This more than doubles the number of known high mass X-ray binary systems in this nearby galaxy. The spatial distribution of the new candidates is similar to that of the already identified Be/X-ray binaries with a strong concentration along the SMC main body and some systems in the eastern wing. The new candidates contribute mainly to the low-luminosity end of the X-ray luminosity distribution of Be/X-ray binaries. A comparison with the luminosity distribution in the Milky Way reveals no significant differences at the high-luminosity end and the large number of low-luminosity systems in the SMC suggests that many such systems may still be undetected in the Galaxy. The overall ratio of known Be to OB supergiant X-ray binaries in the SMC is an order of magnitude larger than in the Galaxy, however, might show spatial variations. While in the eastern wing the ratio is comparable to that in the Galaxy no supergiant X-ray binary is currently known in the main body of the SMC. Possible explanations include a different star formation history over the last 15 My.

Accepted by: Astronomy & Astrophysics For preprints, contact fwh@mpe.mpg.de

Also available from the URL ftp://ftp.xray.mpe.mpg.de/people/fwh/SMC BeX/

Multiwavelength analyses of the extraordinary nova LMC 1991

Greg J. Schwarz^{1,6}, S. N. Shore^{2}, S. Starrfield^{1}, Peter H. Hauschildt^{3}, M. Della Valle^{4},

and E. Baron^{5}

1 Dept. of Physics and Astronomy, Arizona State University, Tempe, AZ 85287-1504, USA 2 Dept. of Physics and Astronomy, Indiana University South Bend, Indiana University South Bend, IN 46634-7111 USA 3 Dept. of Physics and Astronomy, University of Georgia, Athens, GA 30602-2451, USA 4 Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, 50125, Firenze, Italy 5 Dept. of Physics and Astronomy, University of Oklahoma, 440 W. Brooks, Rm 131, Norman, OK 73019-0225, USA 6 Present address: Steward Observatory, University of Arizona, Tucson, AZ 85721, USA

LMC 1991 was a very fast, classical nova and the brightest nova ever observed in the Large Magellanic Cloud. It was extensively observed during both its early optically thick and its nebular evolution in the optical and UV wavelength regions. We have successfully fit all the optically thick spectra using a grid of spherically symmetric, non-LTE, line-blanketed, expanding synthetic spectra created with the model atmosphere code PHOENIX. The emission lines of the nebular spectra have been fit using an optimization technique for the emission line luminosities predicted by the photoionization code CLOUDY. Our analyses show the following: the bolometric luminosity was super-Eddington before visual maximum and reached L_{Bol}^{max} ∼ 6×10^{5} L_{⊙}, the ejected mass was ∼ 3×10^{−4}M_{⊙}, and nuclear burning on the white dwarf ceased after ∼ 100 days. The elemental abundances (by number) with respect to solar of the ejecta were: He/H = 0.8±0.2, C/H = 5_{−2}^{+2.5}, N/H = 85_{−17}^{+22}, O/H = 6.5_{−1.5}^{+2.5}, and all other elements ∼ 0.1. These abundances were determined from both the optically thick and nebular analyses. The extreme luminosity, high ejected mass, rapid turn off time, and low metallicity (except for enhanced CNO) represent the extreme values of observational nova parameters and thus present interesting challenges for understanding the nature of the outburst of LMC 1991.

Submitted to: Monthly Notices of the Royal Astronomical Society For preprints, contact Greg Schwarz, Steward Observatory, University of Arizona, Tucson, AZ 85721

gschwarz@as.arizona.edu

The MACHO Project LMC Variable Star Inventory. IX. Frequency Analysis of the First Overtone RR Lyrae Stars and the Indication for Nonradial Pulsations

Alcock, C.^{1}, Allsman, R.^{2}, Alves, D. R.^{3}, Axelrod, T.^{2}, Becker, A.^{4}, Bennett, D.^{12}, Clement, C.^{15}, Cook, K.H.^{1}, Drake, A.,^{2}, Freeman, K.^{2}, Geha, M.^{1}, Griest, K.^{5}, Kovács, G.^{13}, Kurtz, D.W.^{14}, Lehner, M.^{11}, Marshall, S.^{1}, Minniti, D.^{6}, Nelson, C.^{1,7}, Peterson,

B.^{2}, Popowski, P.^{1}, Pratt, M.^{4}, Quinn, P.^{8}, Rodgers, A.^{16}, Rowe, J.^{15}, Stubbs, C.^{4}, Sutherland, W.^{9}, Tomaney, A.^{4}, Vandehei, T.^{5} and Welch, D.L.^{10}

1Institute for Geophysics and Planetary Physics, Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 2Mount Stromlo and Siding Spring Observatory, Australian National University, Weston Creek, ACT 0200, Australia 3Space Telescope Science Institute, Baltimore, MD 21218, USA 4Department of Astronomy, University of Washington, Seattle, WA 98195, USA 5Department of Physics, University of California, San Diego, CA 92093, USA 6Department Astronomia, Universidad Catolica, Chile 7Department of Physics, University of California, Berkeley, CA 94720, USA 8European Southern Observatory, Garching, D-85748, Germany 9Department of Physics, Oxford University, OX1 3RH, England 10Department of Physics and Astronomy, McMaster University, Hamilton, L82 4M1, Canada 11Department of Physics, University of Sheffield, S3 7RH, England 12Department of Physics, University of Notre Dame, South Bend, IN 46556, USA 13Konkoly Observatory, P.O. Box 67, H-1525 Budapest, Hungary 14Department of Astronomy, University of Cape Town, Rondebosch 7701, South Africa 15Department of Astronomy, University of Toronto, Toronto, M5S 3H8, Canada 16deceased

More than 1300 variables classified provisionally as first overtone RR Lyrae pulsators in the macho variable star database of the Large Magellanic Cloud (LMC) have been subjected to standard frequency analysis. Based on the remnant power in the prewhitened spectra, we found 70% of the total population to be monoperiodic. The remaining 30% (411 stars) are classified as one of 9 types according to their frequency spectra. Several types of RR Lyrae pulsational behavior are clearly identified here for the first time. Together with the earlier discovered double-mode (fundamental & first overtone) variables this study increased the number of the known double-mode stars in the LMC to 181. During the total 6.5 yr time span of the data, 10% of the stars showed strong period changes. We also discovered two additional types of multifrequency pulsators with low occurrence rates of 2% for each. In the first type there remains one closely spaced component after prewhitening by the main pulsation frequency. In the second type the number of remnant components is two, they are also closely spaced, and, in addition, they are symmetric in their frequency spacing relative to the central component. This latter type of variables is associated with their relatives among the fundamental pulsators, known as Blazhko variables. Their high frequency (≈ 20%) among the fundamental mode variables versus the low occurrence rate of their first overtone counterparts makes it more difficult to explain Blazhko phenomenon by any theory depending mainly on the role of aspect angle or magnetic field. Current theoretical models invoke nonradial pulsation components in these stars.

Accepted by: The Astrophysical Journal For preprints, contact kovacs@konkoly.hu Also available from astro-ph/0005361

Abstracts of Non-Refereed Papers

A New Component of the Galaxy as the Origin of the LMC Microlensing Events

E. Gates^{1,2} and G. Gyuk^{3}

1 Adler Planetarium & Astronomy Museum, Chicago, IL 60605, USA 2 Department of Astronomy & Astrophysics, University of Chicago, Chicago, IL 60637, USA 3 Physics Department, University of California, San Diego, CA 92093, USA

We suggest a new component of the Milky Way that can account for both the optical depth and the event durations implied by microlensing searches targeting the Large Magellanic Cloud. This component, which represents less than 4% of the total dark matter halo mass, consists of mainly old white dwarf stars in a distribution that extends beyond the thin and thick disks, but lies well within the dark matter halo. It is consistent with recent evidence for a significant population of faint white dwarfs detected in a proper motion study of the Hubble Deep Field that cannot be accounted for by stars in the disk or spheroid. Further, it evades all of the current observational constraints that restrict a halo population of white dwarfs.

To appear in: Microlensing 2000: A New Era of Microlensing Astrophysics; ASP Conference Series, J.W. Menzies and P.D. Sackett, eds.

Microlensing searches of dark matter

Esteban Roulet

Depto. de Física, Universidad Nacional de La Plata, CC67, 1900, Argentina

The evolution of the observational results of microlensing towards the LMC and some of the suggested interpretations to account for them are discussed. It is emphasized that the results at present are indicative of a lensing population of white dwarfs, possibly in the spheroid (not dark halo) of the Galaxy, together with the more standard backgrounds of stellar populations in the Magellanic Clouds and in the Galaxy. This is also hinted by dynamical estimates of the spheroid mass and by recent direct searches of old white dwarfs.

To appear in: Proceedings of the Silafae III Conference

Available from the URL http://xxx.lanl.gov/ps/astro-ph/0004342

The Star Formation History of the Local Group

Eva K. Grebel^{1,2,3}

1 University of Washington, Dept. of Astronomy, Box 351580, Seattle, WA 98195, USA 2 Hubble Fellow 3 Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany

The star formation histories of Local Group galaxies are summarized. The three large spirals are discussed individually, whereas the discussion of the Local Group dwarfs concentrates on differences and commonalities. While Local Group galaxies exhibit an amazing diversity in their star formation and enrichment histories even within the same morphological type, they share a number of common global properties. Major determining factors for their evolution are galaxy mass and environmental effects such as ram pressure and tidal stripping. It seems likely that low-mass dwarf irregular galaxies evolve into dwarf spheroidals.

To appear in: Proceedings of the 33rd ESLAB Symp. on “Star Formation from the Small to the Large Scale”, eds. F. Favata, A.A. Kaas, & A. Wilson, ESA SP-445, 2000

For preprints, contact grebel@mpia-hd.mpg.de
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0005296